OpenStudy (anonymous):

What is the sum of the geometric sequence 3, 12, 48, … if there are 8 terms?

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

see sum of geometric series = \[a_1 \times \frac{ r^n - 1 }{ r - 1 }\] where r is common ratio , n is no of terms and a_1 is the first term

OpenStudy (anonymous):

But itsn't it a geometric sequence?

rishavraj (rishavraj):

yup it is a geometric sequence..i mentioned it ...

OpenStudy (anonymous):

Sorry! You wrote series so I got confused. Anyways, I got 65535 as an answer?

rishavraj (rishavraj):

yup :))

OpenStudy (anonymous):

Yay! Will you help with one more?

rishavraj (rishavraj):

yeah :))

OpenStudy (anonymous):

Given the arithmetic sequence an = 6 − 4(n + 2), what is the domain for n?

rishavraj (rishavraj):

see n is no or terms so it can only hav positive values

OpenStudy (anonymous):

so n>1?

OpenStudy (anonymous):

Thanks:)

rishavraj (rishavraj):

\[n \ge 1\]

OpenStudy (anonymous):

yes! Thanks so much:)